GIFmaker8:

Divide and reduce to lowest terms. 2 1/4÷4/5 2 13 / 16 2/5 2 1/4 2 4 / 20

GIFmaker8:

@MrMudd183

GIFmaker8:

@MrMudd183

MrMudd183:

i was afk, let me take a look at it

MrMudd183:

i think it's C. with the math i did *improper fractions and cross multiplying*

jhonyy9:

so like a first step make from 2 1/4 an fraction format without integer 2 1/4 = (2*4 +1)/4 = (8+1)/4 = 9/4 and now there are 9/4 ÷ 4/5 = so now hope you know the way how you divide two fractions - so multiplie the first one by invers of second what mean 9/4 ÷ 4/5 = 9/4 × 5/4 = ? so hope you know the way how you multiplie two fractions hope helped understandably easy